# A Physics-Informed, Deep Double Reservoir Network for Forecasting Boundary Layer Velocity

### v2

@@ Abstract
When a fluid flows over a solid surface, it creates a thin boundary layer where the flow velocity is influenced by the surface through viscosity, and can transition from laminar to turbulent at sufficiently high speeds. Understanding and forecasting the fluid dynamics under these conditions is one of the most challenging scientific problems in fluid dynamics. It is therefore of high interest to formulate models able to capture the nonlinear spatio-temporal velocity structure as well as produce forecasts in a computationally efficient manner. Traditional statistical approaches are limited in their ability to produce timely forecasts of complex, nonlinear spatio-temporal structures which are at the same time able to incorporate the underlying flow physics. In this work, we propose a model to accurately forecast boundary layer velocities with a deep double reservoir computing network which is capable of capturing the complex, nonlinear dynamics of the boundary layer while at the same time incorporating physical constraints via a penalty obtained by a Partial Differential Equation (PDE). Simulation studies on a one-dimensional viscous fluid demonstrate how the proposed model is able to produce accurate forecasts while simultaneously accounting for energy loss. The application focuses on boundary layer data on a water tunnel with a PDE penalty derived from an appropriate simplification of the Navier-Stokes equations, showing forecasts improved by $33.7\%$ and $80.0\%$ in terms of mass conservation and variability of velocity fluctuation, respectfully, against non physics-informed methods.
@@ 1 Introduction
Understanding the dynamics and properties of the fluid boundary layer — the thin region of flow near a solid surface where the flow is influenced by the boundary — requires the incorporation of many processes that occur across a range of temporal (i.e., seconds to hours) and spatial (i.e., centimeters to hundreds of meters) scales (Stull, 1988). Boundary layer dynamics are well-studied but complex, and are highly relevant to a range of applications in both engineering and atmospheric science, including wind turbine energy production (Clifton et al., 2014), air pollution transport and dispersion (Fernando et al., 2010), agriculture (Fernando and Weil, 2010) and urban habitability (Conry et al., 2015).
Boundary layer velocities behave as a nonlinear spatio-temporal process whose dynamics can be described by the Navier-Stokes equations, which encode fundamental physical constraints based on conservation of mass and momentum. Recent progresses in machine learning have prompted a different, data-driven approach to modeling and forecasting (see, e.g., Huang and Chalabi (1995); Lei et al. (2009); Ak et al. (2016); Khosravi et al. (2018); Demolli et al. (2019); Huang et al. (2021)). Such methods are predicated on the use of highly non-parametric constructs such as deep Neural Networks (NN) for modeling the space-time dynamics, with increasingly accurate results. Despite these successful implementations, this class of models brings new challenges and questions related to the forecasts.
@@
In this work, we aim at tackling the challenges of practicality and interpretability of forecasting boundary layer velocity by means of a new model that leverages all of the aforementioned features and merges them in a single framework. Specifically, we define a hybrid DESN and DLSM model (deep double reservoir network) which is able to leverage both continuous and binary input. We rely on a physics-informed penalized minimization with a PDE penalty, and we present a computationally affordable algorithm to achieve inference. Finally, we propose a calibration step to achieve proper uncertainty quantification in the forecasts. In the simulation study, we use a simplified one-dimensional viscous fluid, and we show how the proposed approach produces improved forecasts against both statistical and machine learning methods for space-time data. Additionally, our method is able to predict a consistent energy loss in the system, a feature that a data-driven models cannot achieve. Finally, in the application we consider high resolution observations of a turbulent boundary layer in a water tunnel, and penalize its forecast with a simplified Navier-Stokes equation. This type of laboratory experiment provides a scale model of the turbulent motions observed in the atmosphere, as the motion of water can be used to replicate the motions of air given the similar behavior at different length and time scales (dynamic similarity). The resulting model produces accurate forecasts which are also physically consistent and well calibrated. From a practical point of view, this work is meant to serve as a foundation for producing physically-informed and realistic forecasts of atmospheric data and simulations.
@@ 2 Data
We consider experimentally obtained velocity data of a turbulent boundary layer from time-resolved particle image velocimetry (PIV, de Kat and Ganapathisubramani (2015); Fiscaletti et al. (2018)), across a spatial field for over 2,500 time steps at frequency of 1,000 Hz, so that the collection window is 2.5s. Based on the experimental conditions, the underlying space-time process can be described accurately by the incompressible Navier-Stokes equations, which will be used as physics-based penalization in Section 3.4. These data were originally obtained to provide high temporal and spatial resolution of the boundary layer, to better understand the spectral energy characteristics of the near-wall turbulence (de Kat and Ganapathisubramani, 2015; Fiscaletti et al., 2018). Under the principle of dynamic similarity, in which flows are found to be self-similar at some length and time scales, these data also serve as a scale model of the types of motions one would observe in, for example, the atmosphere over minutes or hours. The data are collected on a Cartesian grid of approximate size $0.19$m×$\times$$0.05$m, where the points are equally spaced with horizontal ($x$) and vertical ($y$) resolution of 0.00037m; overall, there are over 62,000 spatial locations with two-dimensional velocity data. The bulk flow characteristics are provided as part of the experiment and include the boundary layer thickness, $\delta=0.1m$; friction velocity, $u_{\tau}=0.027ms^{-1}$; and Reynolds number, $Re_{\tau}=2,700$. In this controlled experiment we use each of these parameters when implementing the physics-informed penalization using a time-averaged version of the Navier-Stokes equations in Section 5. The data were filtered to a frequency of 200 Hz (aggregated in time by a factor of 5$\times$), so that the total number of time points is now $T=500$. Additionally, we subsampled $N=100$ points on a equally spaced diagonal grid from the original spatial locations as shown in Figure 1. The choice of a diagonal grid allows us to retain the important information in the vertical ($y$) direction since the vertical spacing between locations is small ($<$0.004m or 8% of the entire vertical spacing) while also allowing us to offset some possible minor patterns existing in the $x$ direction. The forecast for the $N$ locations is interpolated with thin plate splines (Green and Silverman, 1994) to obtain the remaining spatial domain. Figure 1 shows the entire spatial domain for both the initial time point (panel A) and at $t=0.5s$ (panel B). The black points in these panels represent the $N$ subsampled locations, which will be used for point prediction, physical constraints, and uncertainty quantification in Section 5. Throughout this work, we denote the velocity field by $\mathbf{Y}_{t}=(Y_{t}(\mathbf{s}_{1}),\ldots,Y_{t}(\mathbf{s}_{N}))^{\top}$ at locations $\mathbf{s}_{1},\ldots,\mathbf{s}_{N}$ and times $t=1,\ldots,T$. In the simulation study we employ $\mathbf{Y}_{t}$ directly, while in the application we focused on a time averaged version.
@@ 3.2 Deep Liquid State Machines
Traditional machine learning models such as feedforward NNs compute and transmit continuous valued signals throughout the network. Unlike these, spiking NNs rely on discrete spike streams (binary sequences) to process information, and were formulated to more closely mimic natural biological NNs (Maass, 1997). In this work, we formulate a dynamic version of a spiking NN, a DLSM (Maass et al., 2002), which discretizes the input through a leaky integrate-and-fire model (LIF, Burkitt (2006); Zhang et al. (2018)), which we detail in Section 3.2.1. In Section 3.2.2 we provide the mathematical description of the DLSM. The notation in Section 3.2.1 will be used exclusively for introducing the LIF model, whereas in Section 3.2.2 a slightly different notation will be used, in order to align it with the DESN in the previous Section.
@@ 4 Simulation Study
Reservoir computing methods such as DESN have been shown to produce computationally affordable forecasts owing to the use of random sparse matrices (McDermott and Wikle, 2017; Huang et al., 2021; Bonas and Castruccio, 2023; Yoo and Wikle, 2023), and the PIDDRN is no exception. Even with the addition of the DLSM and the physics-informed loss function, inference can still be achieved in a reasonable time. For this simulation study, we were able to produce 500 ensemble forecasts using the PIDDRN in parallel using only 24 cores at 2.50Ghz frequency in less than 12 hours. This approach could easily be up-scaled by increasing the number of cores.
@@ 4.1 Viscous Burgers’ Equation
where $Y_{t}(x)$ is the fluid velocity at some location $x$ and time point $t$. The Burgers’ equation operates in continuous time, but as assume we observe data at equally spaced times using a discretized version of it. Indeed, we simulate data at $N=30$ uniformly distributed spatial locations on $x\in[-1,1]$ across 500 equally spaced time steps. We solve the equation using a spatial discretization of 1/15 and temporal discretization of 3/1000 (so a total time of 1.5), a choice which allows for a numerically stable solution to the PDE. We specify two different initial conditions: a sinusoid $Y_{0}(x)=-\sin(\pi x)$ and a Gaussian curve $Y_{0}(x)=e^{-x^{2}}$. Additionally, we consider viscosity values of $\nu=\left\{10^{-s};s=1,\dots,4\right\}$, for a total of 8 different simulation setups. The simulated data with $Y_{0}(x)=-\sin(\pi x)$ and $\nu=0.01$ can be seen in Figure 2A, as well as the velocity $Y_{t}(x)$ for all $x$ at times $t=0$, $t=100$, and $t=500$ in Figure 2B. Each model was trained for each of the simulations using the first $T=450$ time steps and was tested using the next $\tau=50$ steps. We produced forecasts using 1000 ensembles, each representing an independent draw for the weight matrices $\mathbf{W}_{d}$ and $\mathbf{W}^{\text{in}}_{d}$ in equations (1f), (1g) and (3). The grids used with the validation approach to optimize the hyper-parameter space are shown in the supplementary material in Table S1. In Section S.2.1 of the supplementary material we further show the ability of the proposed PIDDRN model to produce accurate long lead forecasts on a generalized Burgers’ equation with an external forcing (Büyükaşık and Pashaev, 2013).
@@ 4.2 Point Forecasts and Uncertainty Quantification
The forecasting skills are assessed using DESN as a reference model, and the relative performance of every aforementioned model is computed with respect to MSEpred+${}_{\text{pred}+\text{phys}}$ in equation (5). The results are shown in Table 1, with the median performance across all viscosity values and the interquartile range (IQR) in parenthesis. It can easily be seen how the PIDDRN dramatically outperforms all other models. More specifically, the PIDDRN model returns an average improvement of MSEpred+${}_{\text{pred}+\text{phys}}$ for the sinusoidal initial condition of 48.44% (with IQR $115.22\%$) and for the Gaussian initial condition of 62.94% ($12.37\%$). The PIDDRN also shows a dramatically improved prediction against ARFIMA, whose negative values indicate worse performance than the reference DESN for both initial conditions. Additionally, the PIDDRN model outperforms its non-physics-informed counterpart, the DDRN, which has a smaller improvement over the standard DESN of 13.52% (9.00%) and 38.07% (13.78%) for the two initial conditions. Finally, the PIDDRN model strongly outperforms the physics-informed version of the reference DESN, the PIDESN, which itself only obtained a modest improvement of 26.36% (30.37%) and 15.41% (14.78%).
@@
PIDDRN 48.44 (115.22) 62.94 (12.37)
DDRN 13.52 (9.00) 38.07 (13.78)
PIDESN 26.36 (30.37) 15.41 (14.78)
DESN
ARFIMA -164.65 (354.40) -61.88 (40.04)
Long Short-Term Memory 13.61 (143.12) -6.38 (60.23)
Persistence (Naïve) 15.03 (30.37) 56.22 (10.65)
Table 1: Percent improvement of forecasting performance for each model versus the DESN in terms of the MSEpred+${}_{\text{pred}+\text{phys}}$ metric described in equation (5). A median across all simulated datasets is shown with the IQR in parenthesis.
@@ 5.1 Point Forecasts
We will show the forecasting performance of the PIDDRN model in terms of the physics-based loss for the 100 subsampled locations referenced in Section 2 against the other similar reservoir computing models, the time series approach ARFIMA, the machine learning approach LSTM and the naïve forecasting approach persistence in terms of the time average of the forecast $\bar{Y}$. Similar to Table 1, Table 3 shows the relative improvement in terms of the forecasts on the test data using the newly defined loss function against the standard DESN model. As it is shown, this table illustrates how the PIDDRN again outperforms each of the other methods presented. In fact, the PIDDRN model produces a median 49.96% (with IQR of 23.23%) improvement over the DESN for this application, a result further illustrated in Figure 4. Additionally, the PIDDRN model has a more significant improvement over the next best models, the DDRN and PIDESN models, where they only produce an improvement of 1.67% (9.79%) and 9.57% (33.76%), respectively. As in the case of the simulation study, the DDRN does show improvement against the DESN model, yet it still under-performs relative its physics-informed counterpart, the PIDDRN. Finally, the time series model ARFIMA and machine learning model LSTM are the only models in this instance to produce deteriorating results in terms of the MSE, for they produce results which are 10.99% and 3.22% worse than that of the DESN, respectively.
@@
Forecasting Method Relative Improvement
PIDDRN 49.96 (23.23)
DDRN 1.67 (9.79)
PIDESN 9.57 (33.76)
DESN
ARFIMA -10.99 (44.85)
Long Short-Term Memory -3.22 (65.33)
Persistence (Naïve) -0.66 (54.12)
Table 3: Percent improvement of forecasting performance for each model versus the DESN in terms of the median MSEpred+${}_{\text{pred}+\text{phys}}$ (with IQR) across locations described in equation (5) for the water field application data.

### v3

@@ Abstract
When a fluid flows over a solid surface, it creates a thin boundary layer where the flow velocity is influenced by the surface through viscosity, and can transition from laminar to turbulent at sufficiently high speeds. Understanding and forecasting the fluid dynamics under these conditions is one of the most challenging scientific problems in fluid dynamics. It is therefore of high interest to formulate models able to capture the nonlinear spatio-temporal velocity structure as well as produce forecasts in a computationally efficient manner. Traditional statistical approaches are limited in their ability to produce timely forecasts of complex, nonlinear spatio-temporal structures which are at the same time able to incorporate the underlying flow physics. In this work, we propose a model to accurately forecast boundary layer velocities with a deep double reservoir computing network which is capable of capturing the complex, nonlinear dynamics of the boundary layer while at the same time incorporating physical constraints via a penalty obtained by a Partial Differential Equation (PDE). Simulation studies on a one-dimensional viscous fluid demonstrate how the proposed model is able to produce accurate forecasts while simultaneously accounting for energy loss. The application focuses on boundary layer data in a water tunnel with a PDE penalty derived from an appropriate simplification of the Navier-Stokes equations, showing improved forecasting by the proposed approach in terms of mass conservation and variability of velocity fluctuation against non-physics-informed methods.
@@ 1 Introduction
Understanding the dynamics and properties of the fluid boundary layer — the thin region of flow near a solid surface where the flow is influenced by the boundary — requires the incorporation of many processes that occur across a range of temporal (i.e., seconds to hours) and spatial (i.e., centimeters to hundreds of meters) scales (Stull, 1988). These boundary layer dynamics are well-studied in the field of fluid mechanics but are complex and highly relevant to a range of applications in both engineering and atmospheric science, including wind turbine energy production (Clifton et al., 2014), air pollution transport and dispersion (Fernando et al., 2010), agriculture (Fernando and Weil, 2010) and urban habitability (Conry et al., 2015).
In the context of this work, “fluid” refers broadly to any medium (i.e., liquid, gas, or plasma) that is governed by a continuum description of motion and response to forcing. As such, the velocities of interest behave as a nonlinear spatio-temporal process whose dynamics can be described by the Navier-Stokes equations, which encode fundamental physical constraints based on conservation of mass and momentum. Recent progresses in machine learning have prompted a different, data-driven approach to modeling and forecasting (see, e.g., Huang and Chalabi (1995); Lei et al. (2009); Ak et al. (2016); Khosravi et al. (2018); Demolli et al. (2019); Huang et al. (2021)). Such methods are predicated on the use of highly non-parametric constructs such as deep Neural Networks (NN) for modeling the space-time dynamics, with increasingly accurate results. Despite these successful implementations, this class of models brings new challenges and questions related to the forecasts.
@@
In this work, we aim at tackling the challenges of practicality and interpretability of forecasting boundary layer velocity by means of a new model that leverages all of the aforementioned features and merges them in a single framework. Specifically, we define a hybrid DESN and DLSM model (deep double reservoir network) which is able to leverage both continuous and binary input. We rely on a physics-informed penalized minimization with a PDE penalty, and we present a computationally affordable algorithm to achieve inference. Finally, we propose a calibration step to achieve proper uncertainty quantification in the forecasts. In the simulation study, we use a simplified one-dimensional viscous fluid, and we show how the proposed approach produces improved forecasts against both statistical and machine learning methods for space-time data. Additionally, our method is able to predict a consistent energy loss in the system, a feature that a data-driven models cannot achieve. Finally, in the application we consider high resolution laboratory observations of a turbulent boundary layer, and penalize its forecast with a simplified form of the Navier-Stokes equations. The resulting model produces accurate forecasts which are also physically consistent and well calibrated. From a practical point of view, this work is meant to serve as a foundation for producing physically-informed and realistic forecasts of atmospheric data and simulations.
@@ 2 Data
We consider experimentally obtained velocity data of a turbulent boundary layer from time-resolved particle image velocimetry (PIV, de Kat and Ganapathisubramani (2015); Fiscaletti et al. (2018)), across a spatial field for over 2,500 time steps at frequency of 1,000 Hz, so that the collection window is 2.5s. The data are collected from an experimental water tunnel where the underlying space-time process can be described by the incompressible Navier-Stokes equations, which will be used as physics-based penalization in Section 3.4. This type of laboratory experiment — where the working fluid is water instead of air — is a prototypical example of the concept of dynamic similarity in fluid mechanics. According to this principle, it is the relative strengths of the dominant physical mechanisms which matters, and not the individual, dimensional fluid properties. A common example of this is the dimensionless Reynolds number, which compares the strength of fluid inertia to the viscous forces. Flows of air or water with the same Reynolds number are considered dynamically similar, enabling the use of scale models for studying flows at different dimensional scales (e.g., wind tunnel studies of large-scale flows). In the current context, the water tunnel experiments represent a scale model of the atmospheric boundary layer, where motions on the order of seconds in the tunnel represent motions on the order of minutes or hours in the atmospheric boundary layer.
@@
These data were originally obtained to provide high temporal and spatial resolution of the boundary layer, to better understand the spectral energy characteristics of the near-wall turbulence (de Kat and Ganapathisubramani, 2015; Fiscaletti et al., 2018). Under the principle of dynamic similarity, in which flows are found to be self-similar at some length and time scales, these data also serve as a scale model of the types of motions one would observe in, for example, the atmosphere over minutes or hours. The data are collected on a Cartesian grid of approximate size $0.19$m×$\times$$0.05$m, where the points are equally spaced with horizontal ($x$) and vertical ($y$) resolution of 0.00037m; overall, there are over 62,000 spatial locations with two-dimensional velocity data. The bulk flow characteristics are provided as part of the experiment and include the boundary layer thickness, $\delta=0.1m$; friction velocity, $u_{\tau}=0.027ms^{-1}$; and Reynolds number, $Re_{\tau}=2,700$. This Reynolds number is large, allowing the tunnel experiments to physically represent a wide range of flows (including the atmospheric boundary layer) where fluid inertia strongly dominates over the viscous forces trying to dissipate kinetic energy.
In this controlled water field experiment we use each of these parameters when implementing the physics-informed penalization using a time-averaged version of the Navier-Stokes equations in Section 5. The data were filtered to a frequency of 200 Hz (aggregated in time by a factor of 5$\times$), so that the total number of time points is now $T=500$. Additionally, we subsampled $N=100$ points on a equally spaced diagonal grid from the original spatial locations as shown in Figure 1. The choice of a diagonal grid allows us to retain the important information in the vertical ($y$) direction since the vertical spacing between locations is small ($<$0.004m or 8% of the entire vertical spacing) while also allowing us to offset some possible minor patterns existing in the $x$ direction. The forecast for the $N$ locations is interpolated with thin plate splines (Green and Silverman, 1994) to obtain the remaining spatial domain. Figure 1 shows the entire spatial domain for both the initial time point (panel A) and at $t=0.5s$ (panel B). The black points in these panels represent the $N$ subsampled locations, which will be used for point prediction, physical constraints, and uncertainty quantification in Section 5. Throughout this work, we denote the velocity field by $\mathbf{Y}_{t}=(Y_{t}(\mathbf{s}_{1}),\ldots,Y_{t}(\mathbf{s}_{N}))^{\top}$ at locations $\mathbf{s}_{1},\ldots,\mathbf{s}_{N}$ and times $t=1,\ldots,T$. In the simulation study we employ $\mathbf{Y}_{t}$ directly, while in the application we focused on a time averaged version.
@@ 3.2 Deep Liquid State Machines
Traditional machine learning models such as feedforward NNs compute and transmit continuous valued signals throughout the network. Unlike these, spiking NNs rely on discrete spike streams (binary sequences) to process information, and were formulated to more closely mimic natural biological NNs (Maass, 1997). In this work, we formulate a dynamic version of a spiking NN, a DLSM (Maass et al., 2002), which discretizes the input through a leaky integrate-and-fire model (LIF, Burkitt (2006); Zhang et al. (2018)), which we detail in Section 3.2.1. In Section 3.2.2 we provide the mathematical description of the DLSM. The notation in Section 3.2.1 will be used exclusively for introducing the LIF model, whereas in Section 3.2.2 a slightly different notation will be used, in order to align it with the DESN in the previous Section. Additional details about the heuristic behind this model can be found in the supplementary material.
@@ 4 Simulation Study
Reservoir computing methods such as DESN have been shown to produce computationally affordable forecasts owing to the use of random sparse matrices (McDermott and Wikle, 2017; Huang et al., 2021; Bonas and Castruccio, 2023; Yoo and Wikle, 2023), and the PIDDRN is no exception. Even with the addition of the DLSM and the physics-informed loss function, inference can still be achieved in a reasonable time. For this simulation study, we were able to produce 500 ensemble forecasts using the PIDDRN in parallel using only 24 cores at 2.50Ghz frequency in less than 12 hours. This approach could easily be up-scaled by increasing the number of cores. The computational time could also be decreased by using GPUs instead of CPUs.
@@ 4.1 Viscous Burgers’ Equation
where $Y_{t}(x)$ is the fluid velocity at some location $x$ and time point $t$. The Burgers’ equation operates in continuous time, but as assume we observe data at equally spaced times using a discretized version of it. Indeed, we simulate data at $N=30$ uniformly distributed spatial locations on $x\in[-1,1]$ across 500 equally spaced time steps. We solve the equation using a spatial discretization of 1/15 and temporal discretization of 3/1000 (so a total time of 1.5), a choice which allows for a numerically stable solution to the PDE. We specify two different initial conditions: a sinusoid $Y_{0}(x)=-\sin(\pi x)$ and a Gaussian curve $Y_{0}(x)=e^{-x^{2}}$. Additionally, we consider viscosity values of $\nu=\left\{10^{-s};s=1,\dots,4\right\}$, for a total of 8 different simulation setups. The simulated data with $Y_{0}(x)=-\sin(\pi x)$ and $\nu=0.01$ can be seen in Figure 2A, as well as the velocity $Y_{t}(x)$ for all $x$ at times $t=0$, $t=100$, and $t=500$ in Figure 2B. Each model was trained for each of the simulations using the first $T=450$ time steps and was tested using the next $\tau=50$ steps. We produced forecasts using 1000 ensembles, each representing an independent draw for the weight matrices $\mathbf{W}_{d}$ and $\mathbf{W}^{\text{in}}_{d}$ in equations (1f), (1g) and (3). The grids used with the validation approach to optimize the hyper-parameter space are shown in the supplementary material in Table S4. In the supplementary material we further show the ability of the proposed PIDDRN model to produce accurate long lead forecasts on a generalized Burgers’ equation with an external forcing (Büyükaşık and Pashaev, 2013).
@@ 4.2 Point Forecasts and Uncertainty Quantification
The forecasting skills are assessed in terms of the MSEpred+${}_{\text{pred}+\text{phys}}$ metric described in equation (5) and the results are shown in Table 1, with the median performance across all viscosity values and the interquartile range (IQR) in parenthesis. It can easily be seen how the PIDDRN dramatically outperforms all other models. More specifically, the PIDDRN model returns a median MSEpred+${}_{\text{pred}+\text{phys}}$ for the sinusoidal initial condition of 0.04 (with IQR 0.002) and for the Gaussian initial condition of 0.08 (0.07). The results show an improvement against the DESN model of 89.7% and 63.6% for the sinusoidal and Gaussian initial conditions, respectively. The PIDDRN also shows a dramatically improved prediction against ARFIMA, whose median MSEs of 0.69 (0.99) and 0.36 (0.10) indicates significantly worse performance for both initial conditions. Additionally, the PIDDRN model outperforms its non-physics-informed counterpart, the DDRN, which has a smaller improvement over the standard DESN of returning MSE values of 0.33 (0.61) and 0.14 (0.08) for the two initial conditions, respectively. Finally, the PIDDRN model strongly outperforms the physics-informed version of the reference DESN, the PIDESN, which itself only obtained a modest improvement in MSE, returning values of 0.23 (0.36) and 0.19 (0.05).
@@
PIDDRN 0.04 (0.002) 0.08 (0.07)
DDRN 0.33 (0.61) 0.14 (0.08)
PIDESN 0.23 (0.36) 0.19 (0.05)
DESN 0.39 (0.70) 0.22 (0.09)
ARFIMA 0.69 (0.99) 0.36 (0.10)
Long Short-Term Memory 0.33 (0.57) 0.20 (0.05)
Persistence (Naïve) 0.36 (0.70) 0.10 (0.06)
Table 1: Forecasting performance for each model in terms of the MSEpred+${}_{\text{pred}+\text{phys}}$ metric described in equation (5). A median across all simulated datasets (choices of $\nu$) for each initial condition is shown with the IQR in parenthesis.
@@ 5.1 Point Forecasts
We will show the forecasting performance of the PIDDRN model in terms of the physics-based loss for the 100 subsampled locations referenced in Section 2 against the other similar reservoir computing models, the time series approach ARFIMA, the machine learning approach LSTM and the naïve forecasting approach persistence in terms of the time average of the forecast $\bar{Y}$. Similar to Table 1, Table 3 shows the MSEpred+${}_{\text{pred}+\text{phys}}$ metric for each model for the forecast on the test data. As it is shown, this table illustrates how the PIDDRN again outperforms each of the other methods presented. In fact, the PIDDRN model produces a median MSE of 0.133 (with IQR of 0.046) which is an improvement over the DESN of 50.2% which produces values for the MSE of 0.267 (0.168), a result further illustrated in Figure 4. Additionally, the PIDDRN model has a more significant improvement over the next best models, the DDRN and PIDESN models, where they only produce a MSE of of 0.262 (0.167) and 0.241 (0.130), respectively. As in the case of the simulation study, the DDRN does show improvement against the DESN model, yet it still under-performs relative its physics-informed counterpart, the PIDDRN. Finally, the time series model ARFIMA and machine learning model LSTM are the only models in this instance to produce worse results in terms of the MSE versus the DESN approach, for they produce MSEs of 0.296 (0.121) and 0.275 (0.127), respectively.
@@
Forecasting Method MSEpred+${}_{\text{pred}+\text{phys}}$
PIDDRN 0.133 (0.046)
DDRN 0.262 (0.167)
PIDESN 0.241 (0.130)
DESN 0.267 (0.168)
ARFIMA 0.296 (0.121)
Long Short-Term Memory 0.275 (0.127)
Persistence (Naïve) 0.267 (0.119)
Table 3: Forecasting performance for each model in terms of the median MSEpred+${}_{\text{pred}+\text{phys}}$ (IQR in parenthesis) across the 100 subsampled locations described in Section 2 for the water field application data.